OpenStudy (anonymous):

Question 38 of 165 Solve the system of linear equations below. 2x + y = 5 x + y = 4 x = 1, y = 3 x = 1, y = 5 x = 3, y = 1 x = 9, y = -5

OpenStudy (anonymous):

you can subtract the two equations, then the y term will disappear and then you can find x

OpenStudy (anonymous):

\[\left[\begin{matrix}2 & 1 \\ 1 & 1\end{matrix}\right]\left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}5 \\ 4\end{matrix}\right]\]\[|A| = (2 \times 1) - (1 \times 1 ) = 1\]\[A ^{-1} = \left[\begin{matrix}1 & -1 \\ -1 & 2\end{matrix}\right]\]\[\left[\begin{matrix}1 & -1 \\ -1 & 2\end{matrix}\right]\left[\begin{matrix}5 \\ 4\end{matrix}\right]=\left[\begin{matrix}? \\ ?\end{matrix}\right]\]

OpenStudy (anonymous):

so....its....

OpenStudy (anonymous):

do the matrix multiplication and find out :)

OpenStudy (anonymous):

and that is....

OpenStudy (anonymous):

answer isnt that important its how you get there

OpenStudy (anonymous):

noo im saying whats matrix multiplication

OpenStudy (anonymous):

x=1 y=5

OpenStudy (anonymous):



OpenStudy (anonymous):

im on question 103 now so....